An airplaine in taking off from a field makes a run of 2300ft and leaves the ground in 15.0s from the start. (a) what its acceleration assumed constant? (b) with what speed does it leaves the ground?

1
Expert's answer
2022-05-15T17:03:12-0400

(a)

"d=v_0t+at^2/2, \; v_0=0"

"a=2d/t^2=2*2300/15.0^2=20.4\:\rm ft/s^2"

(b)

"d=\frac{v_f+v_0}{2}t, \; v_0=0"

"v_f=2d/t=2*2300/15.0=307\rm \: ft/s"